OpenStudy (anonymous):

Help please!!! medals will be given :) The function below shows the number of car owners f(t), in thousand, in a city in different years t: f(t) = 1.1t2 - 2.5t + 1.5 The average rate of change of f(t) from t = 3 to t = 5 is ______ thousand per year. Numerical Answers Expected!

random231 (random231):

differenciate the function and then put 5 and 3 in place of x

OpenStudy (anonymous):

So what would it be? I don't really get it

OpenStudy (anonymous):

@random231

random231 (random231):

rate of change = 2.2t-2.5. when t=5 rate = 2.2*5 - 2.5 when t =3 rate= 2.2*3-2.5 so avg = ((2.2*5-2.5)+(2.2*3-2.5))/2

OpenStudy (anonymous):

So for the avg what do i do?

random231 (random231):

avg = ((2.2*5-2.5)+(2.2*3-2.5))/2

random231 (random231):

calculate this

OpenStudy (anonymous):

So it would be 6.3?

random231 (random231):

yea

OpenStudy (anonymous):

alright thank you so much! Do you mind helping me with one more?

random231 (random231):

ure welcome